OpenStudy (anonymous):

write a slope1/3 that contain inercept the point (-4 -3)

OpenStudy (anonymous):

i dont get it what ur question

OpenStudy (anonymous):

inercept ? it should be intercept

OpenStudy (anonymous):

it is a typo

OpenStudy (anonymous):

it say to graph

OpenStudy (anonymous):

what's the intercept of a certain point? i never heard of that

OpenStudy (anonymous):

i thank (-4,-3)

OpenStudy (amistre64):

a line that goes thru the point with the given slope

OpenStudy (anonymous):

i really dont get this quation it hard

OpenStudy (amistre64):

given a point (Px,Py) and a slope "m" the equation can be written in many different ways but usually as: y-Py=m(x-Px)

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

i dont know if my answer is right

OpenStudy (amistre64):

me either; care to show it to us?

OpenStudy (anonymous):

oooo only line that through that (-4 -3) and slope 1/3 then y= mx+c where m is a slope and c is a constant solve this by subtituting information we know -3=(1/3)-4+c -3=-4/3+c c=-3+4/3 c=-5/3 then the line is y=(1/3)x-(5/3)

OpenStudy (amistre64):

that form is the "slope intercept" yes

OpenStudy (anonymous):

i ha ve another question

OpenStudy (anonymous):

what is slope intercept bof the line (-1 ,1) and( 0,3)